OpenStudy (anonymous):

The masses of two similar objects are 24kg and 81kg respectively. If the surface area of the larger object is 540 cm square, find the surface area of the smaller object.

OpenStudy (anonymous):

(24/81) = (8/27)^(2/3) = (2/3)^2 = (4/9) (4/9) * 540 = 240 ans: 240